# Cut the Deadwood Out: Post-Training Model Purification with Selective Module Substitution

### v1

@@ 4.2.3 Greedy search for purified model $M_{pure}$
Purifying $M_{v}$ by module substitution
Given a backdoored model $M_{v}$ (characterized by both high CACC on a clean dataset and high ASR on a poisoned dataset), we aim to identify the specific modules and layers that, when substituted, allow the model to maintain strong performance on the clean task (i.e., minimizing $\delta_{\text{acc}}$) while eliminating the backdoor features (i.e., maximizing $\delta_{\text{asr}}$).
Our algorithm, as demonstrated in Figure 2 and Algorithm 1, tracks the parameters selected for substitution using a module set ($\mathcal{M}$) and a layer set ($\mathcal{L}$), e.g., if =$\mathcal{M}=\{O,F,P\}$ and =$\mathcal{L}=\{7,8,9\}$, it indicates that the $O$, $F$, and $P$ module parameters in the 7th, 8th, and 9th layers of the victim model $M_{v}$ should be replaced with the corresponding parameters from proxy model $M_{p}$. For simplicity, we introduce a substitution matrix $S\in\{0,1\}^{|\mathcal{M}|\times|\mathcal{L}|}$ to specify which modules in which layer in the victim model $M_{v}$ are selected for substitution. Note that while the shape of $S$ is determined by the initial sizes of sets $\mathcal{M}$ and $\mathcal{L}$, these two sets are iteratively updated throughout the algorithm. Each row of $S$ corresponds to a layer in the layer set $\mathcal{L}$, and each column corresponds to a module in the module set $\mathcal{M}$. The value of $S[l,m]=1$ indicates that the module $m\in\mathcal{M}$ in layer $l\in\mathcal{L}$ in the victim model $M_{v}$ is selected for substitution, and $S[l,m]=0$ otherwise. Given $S$, we can succinctly represent the parameter substitution as $S\odot M_{p}+(1-S)\odot M_{v}$, where $\odot$ denotes the element-wise product.
We start by selecting all modules in $M_{v}$ to be replaced with those of $M_{p}$, i.e., =$\mathcal{M}=\{K,Q,V,O,F,P\}$, =$\mathcal{L}=\{1,\ldots,L\}$, $S=\mathbf{1}_{|\mathcal{M}|\times|\mathcal{L}|}$. At $t$-th iteration, the algorithm considers two types of updates to $S^{t}$: deselecting a module from the current module set $\mathcal{M}$ (affecting all selected layers) or deselecting a layer from the layer set $\mathcal{L}$ (affecting all selected modules). The update that maximizes the score in Equation 1 is applied to $S^{t+1}$. The process continues until one of two conditions is met: (1) further iterations will not yield a better strategy (controlled by stopping patience $T$); or (2) no additional layers or modules can be removed, i.e., $|\mathcal{M}|=|\mathcal{L}|=1$. Finally,the strategy $S$ that yields a purified model $M_{pure}$ with the highest score is returned. The complexity of our algorithm is a practical quadratic $O(|\mathcal{M}|^{2}+|\mathcal{L}|^{2})$.
@@ 5.1 Experimental settings
Datasets.
We evaluate our method on four datasets: SST-2 (Socher et al., 2013), OLID (Zampieri et al., 2019), MNLI (Williams et al., 2018), and AGNEWS (Zhang et al., 2015). These datasets cover text classification and natural language inference (NLI) tasks, including binary and multi-class classification scenarios, and are widely used for evaluating text backdoor attacks and defenses (Qi et al., 2021a; He et al., 2023; Gupta and Krishna, 2023; Arora et al., 2024).
We download the open-source versions of these datasets provided by HuggingFace (Lhoest et al., 2021). Table 1 presents the statistics of the evaluated datasets. To train backdoored models, we construct poisoned datasets by injecting triggers into the training split, while the validation split is used to construct a clean and a poison test set respectively. For MNLI, we use a random subset of 100,000 samples to reduce overhead.
@@
We study defenses against four prominent types of data-poisoning backdoor attacks: (1) BadNets (Kurita et al., 2020), (2) InsertSent (Dai et al., 2019), (3) Learnable Word Substitution (LWS(Qi et al., 2021d), and (4) HiddenKiller (Qi et al., 2021c). The first two correspond to insertion-based methods, while the latter two involve synonym substitution and syntactic paraphrasing approaches, respectively.
Follow Arora et al. (2024), we use rare words {"cf", "mn", "bb", "tq", "mb" } as triggers for BadNets and phrases {"I watched this movie", "no cross, no crown" } for InsertSent. The poison target labels for each dataset are listed in Table 1. While our main experiments focus on 20% poison rate (in line with the literature (Dai et al., 2019; Qi et al., 2021c; Arora et al., 2024)), i.e., 20% of training samples are poisoned, we also evaluate lower poison rates of 10%, 5% in Section B.2.
@@
Defense Baselines.
We choose six renowned defenses as our baselines, including two data-wise detection methods: (1) ONION (Qi et al., 2021a) and (2) Z-Def. (He et al., 2023), and four model-wise purification methods: (3) PURE (Zhao et al., 2024), (4) TIES (Yadav et al., 2023), (5) DARE (Yu et al., 2024), and (6) WAG (Arora et al., 2024).
Among these, ONION and Z-Def. detect and remove triggers from datasets. ONION identifies outlier words (potential triggers) using language models, e.g., GPT-2 (Radford et al., 2019), while Z-Def. detects spurious correlations between tokens and labels. In contrast, (2)-(6) are model purification methods. PURE purifies the victim model using attention head pruning and normalization techniques. TIES, DARE and WAG merge the victim model with other models to sanitize it.
Evaluation Metrics.
As introduced in  Section 3, we adopt Clean Accuracy (CACC) and Attack Success Rate (ASR) to measure the model performance. CACC is evaluated on a clean test set, where both the samples and labels are ground truth. In contrast, ASR is evaluated on a poisoned test set, where triggers are implanted into each sample, and the labels are flipped to the target class.
@@
Table 2: (Partial) Performance of our method on two datasets compared to baselines under various backdoor attacks on the RoBERTa-large model, with each value averaged over three seeds. We highlight the top-2 lowest ASR results in blue cells, and the highest CACC results in bold. Results for other datasets are in Table 8.
Implementation details
Following Arora et al. (2024), we compare all methods on RoBERTa-large (Liu et al., 2019) and BERT-base-uncased (Devlin et al., 2019). All victim models are fine-tuned on poisoned datasets for three epochs using the Adam optimizer with no weight decay (Kingma and Ba, 2015) and a learning rate of $2\times 10^{-5}$. All defense baselines are implemented based on their respective open-source repositories (see Section A.3). The high-confidence proxy datasets ($\mathcal{D}_{clean}$ and $\mathcal{D}_{poison}$) are extracted following Section 4.2.2. To train proxy models for selection, we use IMDB (Zhang et al., 2015) for the SST-2 victim model, Twitter Abusive (Founta et al., 2018) for the OLID, SNLI (Young et al., 2014) for MNLI, and BBCNews (Greene and Cunningham, 2006) for AGNews.
Our method uses hyperparameter $\alpha$ in Equation 1 to control the trade-off between backdoor defense and task utility. By default, $\alpha$ is set to 0.4 but can be adjusted based on priorities: smaller values (e.g., 0.1) for lower ASR and larger values (e.g., 1.0) for higher utility. The stopping patience $T=5$ is default in Algorithm 1. In Section B.3, we justify these choices and show that our method is NOT particularly sensitive to hyperparameters.
@@ 5.2 Main results
Table 2 presents the performance of our method compared to all baselines under various backdoor attacks on the RoBERTa-large model, with scores averaged over three runs using different seeds. Results on other datasets and architectures are provided in Section B.1. Across all datasets and backdoor attacks, our method consistently ranks among the top 2, regarding backdoor removal performance, with minor harm to clean accuracy. In particular, under the two particularly challenging backdoor tasks, LWS and HiddenKiller, GMS demonstrates significant improvements (i.e., at least 25%) over all baselines—for example, for SST-2, our method reduces the ASR on LWS to 9.7%, compared to 58.8% for the following baseline Z-Def. While Z-Def achieves competitive performance with GMS in defending against BadNet and InsertSent, Z-Def is much less effective against LWS on both datasets. This is because Z-Def relies on lexical and syntactic features to detect outliers in the poisoned dataset, whereas LWS attacks subtly replace words with synonyms, effectively bypassing outlier detection. As for PURE, another competitive recent approach, we found its head-pruning step is highly unstable across both tasks and architectures. For example, it performs well on the BERT-base model in Table 9 (following the settings reported in PURE (Zhao et al., 2024)), but performs poorly on most tasks with RoBERTa-large. We discuss the potential reasons for this instability in Section B.1.
Dataset Model CACC $\uparrow$
SST-2 Benign Victim Model 95.9
Proxy Model (IMDB) 88.7
Purified Model (GMS) 95.8
AGNews Benign Victim Model 95.2
Proxy Model (BBCNews) 78.1
Purified Model (GMS) 95.3
Table 3: CACC of benign victim models preserved.
@@
Dataset Proxy Model ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$
@@
[…]
Victim SST-2 BadNet 95.6 100.0 95.3 4.5
Proxy SST-2 HiddenKiller 95.4 7.9
Victim SST-2 HiddenKiller 95.4 96.7 95.9 14.0
Proxy SST-2 BadNet 95.6 17.1
Victim SST-2 BadNet (Different triggers) 95.6 100.0 87.8 6.8
Proxy IMDB 87.7 8.6
Victim SST-2 BadNet (Same triggers) 95.6 100.0 95.8 100.0
Proxy IMDB 90.1 100.0
5.3 Ablation studies
Utility Preservation for Benign Victim Models
While all our previous experiments focus on malicious victim models, it is also important to consider scenarios where the victim model is benign. Does our method compromise the utility of a clean victim model? As shown in Table 3, our results effectively dispel this concern. When applying GMS to a benign model, the CACC remains unaffected.
The effect of different proxy models
@@ 5.3 Ablation studies
Our method relies on a proxy model for substitution, raising the natural question of how sensitive it is to the selection of proxy models. We tested two scenarios: Scenario 1: We evaluated the performance of using different proxy models trained on homologous datasets, such as IMDB, Yelp, and Amazon (Zhang et al., 2015), to purify a victim model trained on SST-2. As shown in Table 4, our method is largely insensitive to the choice of proxy model across all backdoor attacks. Using any proxy model, GMS effectively defends against all tested attacks. Scenario 2: What if the proxy model itself is backdoored? We tested cases where the victim model and the proxy model were trained on the same dataset with different backdoor attacks or on different datasets with the same attacks. From Table 5, we observed that as long as the backdoor in the proxy model is not identical to that in the victim model (i.e., the same backdoor strategy with the same backdoor trigger), our GMS method consistently removes the backdoor.
Performance under different knowledge priors
While our method uses random sampling and outlier detection to extract the proxy datasets $\mathcal{D}_{clean}$ and $\mathcal{D}_{poison}$, we also evaluate the potential performance gain if more accurate prior knowledge were available. For example, we may receive an accurate clean dataset or an accurate label-flipped poison dataset (i.e., dirty label) from reliable users or annotators. As shown in Table 6, with or without prior knowledge of accurate clean or poisoned data, our method consistently purifies the backdoor, achieving an ASR same as that of a benign model (4.1%). With optimal priors, GMS maintains the CACC of the benign model (95.6%), compared to 93.6% in the worst case without such prior knowledge.
Knowledge Prior Merged SST-2 (BadNet) & IMDB
[…]
Random Sampling Pure Clean Outlier Detection Dirty Label
3.2 93.6
4.1 94.9
4.3 92.7
4.3 95.6
Benign SST-2 4.1 95.6
Table 6: Performance of purifying a victim model trained on SST-2 with a proxy model trained on IMDB, using proxy datasets with varying knowledge priors. The backdoor attack strategy is BadNet.
BadNet InsertSent LWS HiddenKiller
[…]
Benign 4.1 95.9 2.2 95.9 12.8 95.9 16.5 95.9
Victim 100.0 96.0 100.0 96.3 98.0 95.4 96.5 95.7
Proxy Model (IMDB) 7.4 89.1 4.3 89.1 10.8 89.1 13.7 89.1
GMS 4.5 91.6 1.9 92.5 9.7 91.7 10.4 91.2
SST-2 SST-2 BadNet Strategy 4.5 91.6 9.8 93.0 22.3 92.8 13.2 90.3
Benign 1.9 95.4 0.5 95.4 0.5 95.4 1.1 95.4
Victim 99.9 95.1 99.6 95.3 99.6 94.5 100.0 95.1
Proxy Model (BBCNews) 1.5 70.2 1.7 70.2 1.8 70.2 3.4 70.2
GMS 2.5 91.0 2.4 92.6 3.2 91.7 6.5 90.4
AGNews AGNews BadNet Strategy 1.4 90.7 1.2 89.8 2.6 90.1 12.0 90.2
Table 7: (Partial) The performance of transferring strategy searched based on BadNet to other attacks. Results for other datasets are in Table 12.
Transferability of substitution strategies
Interestingly, upon examining the module substitution strategies ($S_{best}$) searched by GMS for addressing various backdoor attacks (illustrated with examples in Figure 6 in Section B.5), we observe that, despite differences among the strategies, they exhibit transferability across different backdoor attacks.
In Table 7, for each task dataset, we apply the substitution strategy found on BadNet (replacing the fewest modules and layers as shown in Figure 6) to to determine which modules to substitute in the three other victim models under different attacks, replacing them with modules from their corresponding proxy models, i.e., directly executing Line 24 in Algorithm 1 using $S_{badnet}$. Compared with the complete GMS (i.e., searching for the optimal strategies $S_{best}$ for each attack), the transferred strategy consistently performs comparably. This demonstrates that for defenders who only have access to the victim model and are unaware of the victim dataset, a universal GMS defense strategy exists for each task that can effectively defend against multiple attacks.

### v2

@@ 4.2.3 Greedy search for purified model $M_{pure}$
Purifying $M_{v}$ by module substitution.
Given a victim model $M_{v}$ (likely to has high ACC and high ASR), we next illustrate how to identify the specific modules and layers that, when substituted, allow the model to maintain strong performance on the clean task (i.e., minimizing $\Delta_{\text{acc}}$) while eliminating the backdoor features (i.e., maximizing $\Delta_{\text{asr}}$).
Our algorithm, as demonstrated in Figure˜1 and Algorithm˜1, tracks the parameters selected for substitution using a module set ($\mathcal{M}$) and a layer set ($\mathcal{L}$), e.g., if $\mathcal{M}=\{O,F,P\}$ and $\mathcal{L}=\{7,8,9\}$, it indicates that the $O$, $F$, and $P$ module parameters in the 7th, 8th, and 9th layers of the victim model $M_{v}$ should be replaced with the corresponding parameters from proxy model $M_{p}$. For simplicity, we introduce a substitution matrix $S\in\{0,1\}^{|\mathcal{M}|\times|\mathcal{L}|}$ to specify which modules in which layer in the victim model $M_{v}$ are selected for substitution. Note that while the shape of $S$ is determined by the initial sizes of sets $\mathcal{M}$ and $\mathcal{L}$, these two sets are iteratively updated throughout the algorithm. Each row of $S$ corresponds to a layer in the layer set $\mathcal{L}$, and each column corresponds to a module in the module set $\mathcal{M}$. The value of $S[l,m]=1$ indicates that the module $m\in\mathcal{M}$ in layer $l\in\mathcal{L}$ in the victim model $M_{v}$ is selected for substitution, and $S[l,m]=0$ otherwise. Given $S$, we can succinctly represent the parameter substitution as $S\odot M_{p}+(1-S)\odot M_{v}$, where $\odot$ denotes the element-wise product.
Since the primary goal is backdoor removal, we start by selecting all modules in $M_{v}$ to be replaced with those of $M_{p}$, i.e., $\mathcal{M}=\{K,Q,V,O,F,P\}$, $\mathcal{L}=\{1,\ldots,L\}$, $S=\mathbf{1}_{|\mathcal{M}|\times|\mathcal{L}|}$. At $t$-th iteration, the algorithm considers two types of updates to $S^{t}$: deselecting a module from the current module set $\mathcal{M}$ (affecting all selected layers) or deselecting a layer from the layer set $\mathcal{L}$ (affecting all selected modules). The update that maximizes the score in Equation˜1 is applied to $S^{t+1}$. The process continues until one of two conditions is met: (1) further iterations will not yield a better strategy (controlled by stopping patience $T$); or (2) no additional layers or modules can be removed, i.e., $|\mathcal{M}|=|\mathcal{L}|=1$. Finally, the strategy $S$ that yields a purified model $M_{pure}$ with the highest score is returned. The complexity of our algorithm is a practical quadratic $O(|\mathcal{M}|^{2}+|\mathcal{L}|^{2})$.
@@ 5.1 Experimental settings
Datasets. We evaluate our method on four datasets: SST-2 (Socher et al., 2013), OLID (Zampieri et al., 2019), MNLI (Williams et al., 2018), and AGNEWS (Zhang et al., 2015). These datasets cover text classification and natural language inference (NLI) tasks, including binary and multi-class classification scenarios, and are widely used for evaluating text backdoor attacks and defenses (Qi et al., 2021a; He et al., 2023; Gupta and Krishna, 2023; Arora et al., 2024). We adapt the open-source datasets provided by HuggingFace (Lhoest et al., 2021). Table 1 are the statistics of these datasets.
@@
We study defenses against four prominent types of data-poisoning backdoor attacks: (1) BadNets (Kurita et al., 2020), (2) InsertSent (Dai et al., 2019), (3) Learnable Word Substitution (LWS(Qi et al., 2021c), and (4) HiddenKiller (Qi et al., 2021b). The first two correspond to insertion-based methods, while the latter two involve synonym substitution and syntactic paraphrasing approaches, respectively. A more recent attack, BITE Yan et al. (2023), is discussed separately in Section˜B.3 due to its atypical behavior on benign models.
Follow Arora et al. (2024), we use rare words {"cf", "mn", "bb", "tq", "mb" } as triggers for BadNets and phrases {"I watched this movie", "no cross, no crown" } for InsertSent. The poison target labels for each dataset are listed in Table 1. While our main experiments focus on 20% poison rate (in line with the literature (Dai et al., 2019; Qi et al., 2021b; Arora et al., 2024)), i.e., 20% of training samples are poisoned, we also evaluate lower poison rates of 10%, 5% in Section˜B.4.
@@
Table 2: (Partial) Performance of our method on two datasets compared to baselines under various backdoor attacks on the RoBERTa-large model, with each value averaged over three seeds. We highlight the top-2 lowest ASR results in blue cells, and the highest CACC results in bold. Results for other datasets are in Table˜8.
Defense Baselines.
We choose seven renowned defenses as main baselines, including two data-wise detection methods: (1) ONION (Qi et al., 2021a) and (2) Z-Def. (He et al., 2023), and four model-wise purification methods: (3) PURE (Zhao et al., 2024b), (4) ABL (Li et al., 2021b), (5) TIES (Yadav et al., 2023), (6) DARE (Yu et al., 2024), and (7) WAG (Arora et al., 2024).
Among these, ONION and Z-Def. detect and remove triggers from datasets. ONION identifies outlier words (potential triggers) using language models, e.g., GPT-2 (Radford et al., 2019), while Z-Def. detects spurious correlations between tokens and labels. In contrast, (2)-(7) are model purification methods. PURE purifies the victim model using attention head pruning and normalization techniques, and ABL introduces a robust anti-backdoor training framework that first isolates, then unlearns backdoor associations. TIES, DARE and WAG are mode-merging baselines. We also report three additional baselines in Section˜B.2: two model purification methods, Fine-mixing (Zhang et al., 2022) and Fine-purifying (Zhang et al., 2023), and one data filtering method, SEEP (He et al., 2024), followed by retraining.
Evaluation Metrics.
Following previous literature (Qi et al., 2021a), we adopt Clean Accuracy (CACC) and Attack Success Rate (ASR) to measure utility and defense performance respectively. CACC is evaluated on a clean test set, where both the samples and labels are ground truth. In contrast, ASR is evaluated on a poisoned test set, where triggers are implanted into each sample, and the labels are flipped to the target class.
Implementation details.
Following Arora et al. (2024), we compare all methods on RoBERTa-large (Liu et al., 2019), BERT-base-uncased (Devlin et al., 2019), as well as LLMs including Llama 2 7B (Touvron et al., 2023), Mistral 7B (Jiang et al., 2023), and Qwen 2.5 7B (Team, 2024), with low-rank adaptation LoRA (Hu et al., 2022). All victim models are fine-tuned on poisoned datasets using the Adam optimizer with no weight decay (Kingma and Ba, 2015) and a learning rate of $2\times 10^{-5}$. All defense baselines are implemented based on their open-source repositories (see Section˜A.3). The high-confidence proxy datasets ($\mathcal{D}_{clean}$ and $\mathcal{D}_{poison}$) are extracted following Section˜4.2.2. To train proxy models for selection, we use IMDB (Zhang et al., 2015) for the SST-2 victim model, Twitter Abusive (Founta et al., 2018) for the OLID, SNLI (Young et al., 2014) for MNLI, and BBCNews (Greene and Cunningham, 2006) for AGNews.
Our method uses hyperparameter $\alpha$ in Equation˜1 to control the trade-off between backdoor defense and task utility. By default, $\alpha$ is set to 0.4 but can be adjusted based on priorities: smaller values (e.g., 0.1) for lower ASR and larger values (e.g., 1.0) for higher utility. The stopping patience $T=5$ is default in Algorithm˜1. In Section˜B.5, we justify these choices and show that our method is NOT particularly sensitive to hyperparameters.
@@ 5.2 Main results
Table˜2 presents the performance of our method compared to all baselines under benchmark backdoor attacks on the RoBERTa-large model, with scores averaged over three runs using different seeds. Results on other datasets and architectures are provided in Section˜B.1. Across all datasets and backdoor attacks, our method consistently ranks among the top 2, regarding backdoor removal performance, with minor harm to clean accuracy. In particular, under the two particularly challenging backdoor tasks, LWS and HiddenKiller, GMS demonstrates significant improvements (i.e., at least 25%) over all baselines–for example, for SST-2, our method reduces the ASR on LWS to 9.7%, compared to 58.8% for the following baseline Z-Def. While Z-Def achieves competitive performance with GMS in defending against BadNet and InsertSent, Z-Def is much less effective against LWS on both datasets. We attribute this to Z-Def.’s reliance on lexical and syntactic features to detect outliers in the poisoned dataset, whereas LWS attacks subtly replace words with synonyms, effectively bypassing outlier detection. As for PURE, another competitive recent approach, we found its head-pruning step is highly unstable across both tasks and architectures. For example, it performs well on the BERT-base model in Table˜9 (following the settings reported in PURE (Zhao et al., 2024b)), but performs poorly on most tasks with RoBERTa-large. Unlearning-based methods (ABL Li et al. (2021b)) exhibit similar issues–often degrading CACC more than reducing ASR. Such instability has also been noted in prior works Liu et al. (2018); Wu et al. (2022), which shows that these methods are sensitive to both hyperparameter choices and the quality of the proxy data. Additionally, we observed that all model-merging baselines suffer in defense performance when merging with a single proxy model, aligning with the results in Arora et al. (2024), and tend to prioritize preserving accuracy over removing backdoors. We discuss the potential reasons for this and the instability of PURE in Section˜B.1. We also compare our approach with two earlier model purification baselines, Fine-mixing Zhang et al. (2022) and Fine-purifying Zhang et al. (2023) in Section˜B.2.
Performance on LLMs.
Given the growing popularity of LLMs, we evaluate the performance of our method on Llama-2-7B, Mistral-7B and Qwen-2.5-7B for the SST-2 dataset (Table˜4). Remarkably, in all cases, GMS effectively removes backdoors.
[…]
Hidden Killer 95.4 4.5
BadNet (Diff. Trigger) 87.8 6.9
BadNet (Same Trigger) 95.8 100.0
Victim Model (BadNet) 95.6 100.0
Table 3: Proxy models with implanted backdoors can still effectively purify victim models if the backdoor is not identical in both attack strategy and trigger.
5.3 Ablation studies
We next examine the key property of GMS–robustness across diverse proxy datasets. For completeness, we defer additional results to the appendix: stability under different hyperparameter settings (Section˜B.5), transferability across attacks (Section˜B.12), preservation of clean victim model utility (Section˜B.6), and resilience to re-tuning attacks (Section˜B.10).
@@ 5.3 Ablation studies
[…]
Benign 3.6 96.7 3.5 96.7 14.9 96.7 15.2 96.7
Victim 100.0 97.0 100.0 97.1 98.7 96.1 95.9 96.7
Llama-2 GMS 3.5 91.7 3.6 93.4 12.2 91.5 7.6 91.1
Benign 3.8 97.1 4.5 97.1 7.7 97.1 12.3 97.1
Victim 100.0 95.8 100.0 96.5 98.9 95.5 95.9 96.4
Mistral GMS 4.4 92.1 7.1 92.4 3.8 91.9 5.3 91.9
Benign 5.0 97.0 2.7 97.0 15.1 97.0 15.7 97.0
Victim 100.0 96.3 100.0 92.9 98.5 98.5 96.0 96.2
Qwen-2.5 GMS 6.3 94.5 7.5 94.4 16.4 91.7 15.0 93.0
Table 4: Performance of our method on Llama-2-7b, Mistral-7b and Qwen-2.5-7b on SST-2 dataset, averaged over three seeds.
BadNet InsertSent LWS HiddenKiller
Victim Model Proxy Model ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$
@@
Proxy-model robustness:
@@
GMS retains the robustness property of model-merging defenses: it is resilient to both the specific choice of the proxy model and the benign or malicious nature of the proxy model; i.e., as long as the backdoors in the two models are not identical, the backdoor can be effectively removed. We examine the sensitivity of GMS to the selection of proxy models in two scenarios. Scenario 1: Different proxies. We evaluated the performance of using different proxy models trained on homologous datasets, such as IMDB, Yelp, and Amazon (Zhang et al., 2015), to purify a victim model trained on SST-2. As shown in Table˜5, our method is largely insensitive to the choice of proxy model across all backdoor attacks. Using any proxy model, GMS effectively defends against all tested attacks. Scenario 2: Backdoored proxies. We tested cases where the victim model and the proxy model were trained on the same dataset (SST2) with different backdoor attacks (BadNet and Hidden Killer) or on different datasets (SST2 and IMDB) with the same attacks (BadNet). From Table˜3 and Table˜17, we observed that as long as the backdoor in proxy model is not identical to that in victim model (i.e., the same backdoor strategy with an identical backdoor trigger), GMS consistently mitigates the backdoor. We attribute this to the combination of module-level substitution and disruptive nature of model merging, which breaks the backdoor pathway such that a non-identical proxy cannot restore it.
Sub-proxy datasets GMS
Proxy Clean Proxy Poison CACC $\uparrow$ ASR $\downarrow$
Random Sample Outlier Detection 93.6 3.2
Random Sample Oracle 94.9 4.1
Oracle Outlier Detection 92.7 4.3
Oracle Oracle 95.6 4.3
Benign SST-2 95.6 4.1
Table 6: Performance comparison of different sub-proxy dataset configurations on GMS.
Proxy-data robustness:
Prior pruning- and unlearning-based methods strictly rely on access to clean data or known backdoor triggers Min et al. (2024); Li et al. (2021b); Chen et al. (2022); Li et al. (2023), typically extracted from the training subset. In contrast, GMS operates under more practical settings, e.g., post-processing after suboptimal data detections, where the data received is a mixture of clean and poisoned samples. Our previous experiments used $\mathcal{D}_{clean}$ from random sampling (containing 20% poisoned data) and $\mathcal{D}_{poison}$ from heuristic outlier detection. Table˜7 shows that even when 70% of the “clean” proxy dataset is poisoned, GMS is still effective in purification. In Appendix˜C, we mathematically justify this and discuss the analytical constraints on proxy datasets. These results demonstrate that our method is highly robust to proxy-data construction. Nevertheless, Table˜6 shows that more accurate proxy datasets–such as an oracle clean or label-flipped poison dataset–can further enhance performance, enabling precise module identification and full utility preservation (e.g., CACC of 95.6%). However, even without optimal proxies, our method consistently removes backdoors, achieving benign-level ASR (4.1%) with comparable CACC.
SST-2 AGNews
Ratio $\rho$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$
0.00 93.58 3.38 91.76 2.42
0.30 91.51 3.15 91.59 3.61
0.60 89.11 5.18 88.67 1.42
0.70 89.11 5.18 88.67 1.42
0.80 89.11 5.18 94.59 99.75
0.90 89.11 5.18 94.83 99.82
Table 7: Purification performance as the ratio $\rho$ of poisoned data in extracted $\mathcal{D}_{clean}$ increases.